OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5 over 9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 20 degrees Celsius is......

OpenStudy (anonymous):

Part 1 : C = 5(F-32)/9 9C/5 = F-32 (9C/5) + 32 = F Part 2 : C = 5(20-32)/9 = -5*12/9 = -60/9 = -20/3 = -6.67 degree Celcius